steviej148:

Since YouTube first became available to the public in mid-2005, the rate at which video has been uploaded to the site can be approximated by v(t) = 525,600(0.42t^2 + 2.7t − 1) hours of video per year (0.5 ≤ t ≤ 7), where t is the time in years since the start of 2005.† (a) Find an expression for the total number of hours V(t) of video at time t (starting from zero hours of video at t = 0.5). (b) Use the answer to part (a) to estimate the total number of hours of video uploaded by the start of 2011. (Round your answer to the nearest million hours of video.) _______ million hours of video

Vocaloid:

for part (a) you are given the rate function and the time, to find the # of videos you can take the integral of v(t) = 525,600(0.42t^2 + 2.7t − 1) between t = 0.5 and v = 7 (try factoring out 525,600 from the integral first)

AdiGirl:

yeah thats right

AdiGirl:

thanks for the answer vocaloid, it helped me!

Vocaloid:

still working on part (b) just a moment

AdiGirl:

yea ik

AdiGirl:

hi Stevie!

Vocaloid:

actually I believe part (a) is just setting up the integral and part (b) is just solving it so for part (a) simply write down the integral and for part (b) evaluate it

AdiGirl:

oki

steviej148:

i think I have the integral right but im pretty sure i have to solve for c as well to be able to get the answer to part b and im strggling with that

Vocaloid:

it's a definite integral so there shouldn't be a c value

AdiGirl:

yea

steviej148:

what did you guys get for the integral?

Vocaloid:

the integral of t^2 is (1/3)t^3 the integral of t is (1/2)t^2 integral of 1 is t putting that together integral of 525,600(0.42t^2 + 2.7t − 1) = 525,600(0.42 * (1/3) * t^3 + 2.7(1/2)t^2 - t) evaluate that at 7 and 0.5 then subtract the two values

AdiGirl:

ye thats true

AdiGirl:

so do you guys like anime????? just curious

Vocaloid:

let's please try and stay on topic (there are chat boxes at the bottom for casual conversation)

AdiGirl:

where are the chat boxes?

Vocaloid:

at the bottom of your screen

AdiGirl:

oki saw it

Vocaloid:

@steviej148 you still there? can you try evaluating: A) 525,600(0.42 * (1/3) * t^3 + 2.7(1/2)t^2 - t) for t = 7 and B) 525,600(0.42 * (1/3) * t^3 + 2.7(1/2)t^2 - t) for t = 0.5 then subtracting A ) - B ) ?

steviej148:

yes one second

AdiGirl:

son¡Steve u like anime?

AdiGirl:

sorry meant to say so

AdiGirl:

why is there so many little people on question cove????